OpenStudy (anonymous):

Solve for x: the quantity of x plus 20 over 2 = 3x x = 10 x = 4 x = -14 x = -17

OpenStudy (anonymous):

x + 20/2 = 3x multiple both sides by 2 to get rid of the 2 2(x+20/2) = 3x(2) the 2s cancel each other out on the left side x+20=6x you want to get the x's on the same side so subtract x from each side 20=5x divide both sides by 5 to get x on its own 20/5 = 5x/5 4=x

OpenStudy (anonymous):

Can you help me with like 4 more? xx

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6. 1 4 5 25

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5. 6 -6 All real numbers No solution

OpenStudy (anonymous):

No solution

OpenStudy (anonymous):

Okay, thank you so much!

OpenStudy (anonymous):

No problem :)